OpenStudy (anonymous):

Find the indicated limit, if it exists. limit of f of x as x approaches 2 where f of x equals x plus 3 when x is less than 2 and f of x equals 3 minus x when x is greater than or equal to 2

OpenStudy (anonymous):

where are you stuck?

OpenStudy (anonymous):

I dont know where to start

OpenStudy (anonymous):

I am having a lot of trouble with this lesson. can you walk me through it please?

OpenStudy (anonymous):

\(lim_{x\rightarrow 2}x+3=??\)

OpenStudy (anonymous):

I will draw it

OpenStudy (anonymous):

just replace x =2 into it. so x+3 =2+3 =5, hence lim =5

OpenStudy (anonymous):

can you help me with this one too? User: Colton Stanger In Course: Pre-Calculus V12 ( 3831) Instructor: David DAgostino Exam: 9.04 Continuity at a Point Warning: There is a checkbox at the bottom of the exam form that you MUST check prior to submitting this exam. Failure to do so may cause your work to be lost. Question 1 (Multiple Choice Worth 2 points) Use graphs and tables to find the limit and identify any vertical asymptotes of limit of 1 divided by the quantity x minus 5 as x approaches 5 from the left. 1 ; no vertical asymptotes -∞ ; x = 5 ∞ ; x = -5 -∞ ; x = -5 Question 2 (Multiple Choice Worth 2 points) Find the indicated limit, if it exists. limit of f of x as x approaches 0 where f of x equals 5 x minus 9 when x is less than 0 and the absolute value of the quantity 2 minus x when x is greater than or equal to 0 -9 -7 2 The limit does not exist. Question 3 (Multiple Choice Worth 2 points) Find the indicated limit, if it exists. limit of f of x as x approaches 5 where f of x equals 5 minus x when x is less than 5, 8 when x equals 5, and x plus 3 when x is greater than 5

OpenStudy (anonymous):

for question1, if x =-5 then f(x) is undefined, hence x =-5 is asymptote

OpenStudy (anonymous):

when x =5 just replace into f(x) , that is its limit

OpenStudy (anonymous):

hence the right one is -infinitive, x =-5